# MultiBLiMP 1.0: A Massively Multilingual Benchmark of Linguistic Minimal Pairs

### v1

@@ Abstract
We introduce MultiBLiMP 1.0, a massively multilingual benchmark of linguistic minimal pairs, covering 101  languages, 6 linguistic phenomena and containing more than 125,000 minimal pairs. Our minimal pairs are created using a fully automated pipeline, leveraging the large-scale linguistic resources of Universal Dependencies and UniMorph. MultiBLiMP 1.0 evaluates abilities of LLMs at an unprecedented multilingual scale, and highlights the shortcomings of the current state-of-the-art in modelling low-resource languages. [^1] Code: github.com/jumelet/multiblimp
@@ 1 Introduction
Large Language Models (LLMs) are often trained on highly multilingual corpora, which enable users to interact with them in a wide range of languages (Grattafiori et al., 2024; Üstün et al., 2024). Multilingual evaluation, however, has mostly focused on their functional linguistic competence through tasks requiring world knowledge and language understanding (Singh et al., 2024a, b), while little work has assessed their formal linguistic competence, i.e. “knowledge of rules and statistical regularities of a language” (Mahowald et al., 2024). In multilingual contexts, the latter is commonly approximated intrinsically through perplexity (Chang et al., 2024) or extrinsically through performance in generative tasks such as translation or summarization (Dang et al., 2024a). These approaches do not truly disentangle formal from functional competence. Moreover, they are very coarse-grained and do not inform us on which specific constructions a model does (not) master.
Figure 1: Gemma3-27B accuracy per language on MultiBLiMP 1.0, plotted against language frequency in Common Crawl. MultiBLiMP enables fine-grained evaluation of linguistic abilities on a language-level.
Figure 2: Pipeline of the minimal pair creation procedure of MultiBLiMP 1.0.
@@
To accelerate progress in this direction, we introduce MultiBLiMP 1.0, a massively multilingual benchmark of linguistic minimal pairs covering subject-verb agreement for number, person, and gender; created automatically using two large-scale linguistic resources: Universal Dependencies (Zeman et al., 2024) and UniMorph (Batsuren et al., 2022). MultiBLiMP is not only a benchmark, but also a pipeline for the automatic creation of highly multilingual benchmarks (Fig. 2), which can scale to many more syntactic phenomena. Taking subject-verb agreement as a use case, we present a first version of the benchmark including more than 125,000 minimal pairs across 101  languages. We use this to evaluate 17 LLMs, finding that linguistic competence is strongly driven by model size and language frequency in training data (Fig. 1), is acquired during pre-training, and can deteriorate by post-training.
Besides its practical relevance for LM evaluation, MultiBLiMP enables the study the learnability of specific linguistic constructions cross-lingually in a unified framework, at a much larger scale than currently possible Mueller et al. (2020), with future applications in LM interpretability (Brinkmann et al., 2025) and quantitative typology (Levshina et al., 2023; Baylor et al., 2024).
@@ 2 Background
2.1 Targeted Syntactic Evaluations
On minimal pairs
Evaluations based on minimal pairs have been criticised as potentially not reflecting real-world behaviour if they do not follow the natural frequency distribution of the data, but infrequent verbs remain important to test the systematicity of the models (Newman et al., 2021; Vamvas and Sennrich, 2021). The best way to evaluate instruction-tuned LLMs on minimal pairs and grammaticality has also been debated: Dentella et al. (2023) found GPT3’s abilities lacking using prompting, but Hu et al. (2024) re-examine the data using perplexity scoring and find models and human judgments highly correlated. It has further been argued that binary notions of grammatical acceptability are reductive Weissweiler et al. (2025), and indeed, Hu et al. (2024) also found better correlation by using gradient values for both models and humans, compared to Dentella et al. (2023) who treated both as binary. We do not perform gradient evaluations here as this would require human judgements for all sentences Juzek (2024). However, it is crucial to note that minimal pairs like those used here are immune to this issue, as sentences are not objectively evaluated as grammatical or ungrammatical, but rather compared, and as they differ only in one morphosyntactic feature, they are immune to semantic effects such as those reported by Kauf et al. (2024).
English
@@ 2.1 Targeted Syntactic Evaluations
Minimal pairs to assess subject-verb agreement in LMs were first introduced by Linzen et al. (2016). Specifically, they extracted sentences from Wikipedia and flipped the number of the verb without detecting the subject, finding that models at that time struggled considerably with the distance between the subject and the verb, as well as intervening nouns and other attractors. This work was extended by Marvin and Linzen (2018), who generated sentences with context-free grammars, enabling them to increase the difficulty of agreement across phrases and clauses. The monolingual predecessor to our work is BLiMP (Warstadt et al., 2020), a set of minimal pairs genereated with templates for many linguistic phenomena by sampling from a vocabulary. Hu et al. (2020) and Gauthier et al. (2020) both present similar large-scale collections of syntactic evaluation for English.
Other monolingual benchmarks
@@
Minimal pair benchmarks have recently been developed for various other languages, using different methods. Some use existing annotated sentences from UD, some craft templates, and some others are built by translating an English dataset. For Chinese, CLiMP (Xiang et al., 2021) generate their minimal pairs with templates adapted from BLiMP, while SLING (Song et al., 2022) edit a constituency treebank with templates and then verify the results with human annotators. ZhoBLiMP (Liu et al., 2024) employ a mixture of pairs translated from BLiMP, extracted from syntax journals and textbooks, manipulated with templates. JBliMP (Someya and Oseki, 2023) extract minimal pairs for from journal articles on Japanese syntax. BLiMP-NL Suijkerbuijk et al. (2024) create ten sentence pairs per paradigm by hand for Dutch, as well as 90 more sentences generated by ChatGPT and then hand-checked. Closer to our approach, RuBLiMP (Taktasheva et al., 2024) create pairs from UD-parsed sentences using perturbation rules. Most of these monolingual benchmarks cover agreement as well as other, possibly language-specific, syntactic phenomena.
@@
Four multilingual minimal pair datasets are known to us. The largest, CLAMS (Mueller et al., 2020), creates pairs for English, German, Hebrew, Italian, and Russian for subject-verb agreement by having native speakers translate the sentences from Marvin and Linzen (2018) where applicable. Gulordava et al. (2018) similarly cover English, Hebrew, Italian, and Russian for subject-verb agreement with sentences from UD treebanks, in which content words are swapped out with others of the same POS to test the models’ reliance on semantics. For low-resource languages, LINDSEA (Leong et al., 2023) present minimal pairs for Tamil and Indonesian which are notably evaluated using prompting with GPT3.5 and GPT-4. Kryvosheieva and Levy (2024) create minimal pairs specifically for Swahili noun class agreement, Hindi split ergativity and Basque verb agreement, each chosen for their complexity.
We see MultiBLiMP 1.0 being complementary to all these efforts. While language-specific benchmarks can provide tremendous depth on the syntactic system of a language, it will prove challenging to scale such efforts to a wide variety of languages. Our approach, on the other hand, focuses on ‘wide coverage’ and can therefore be of use for both the evaluation of highly multilingual LLMs and for quantitative typological studies.
@@ 2.2 Representation Sharing in LLMs
A branch of work has evaluated the degree to which multilingual models make use of shared syntactic representations across languages. Early studies analyzed mBERT (Devlin et al., 2019) and smaller-scale LSTM models trained on a few languages, finding mixed evidence of sharing (Pires et al., 2019; Mueller et al., 2020; Dhar and Bisazza, 2021; Chi et al., 2020). Focusing on modern-scale LLMs, Wendler et al. (2024) find that Llama2 models (Touvron et al., 2023) trained on English-dominated corpora use English to some degree as an internal pivot language to solve tasks in other languages. Brinkmann et al. (2025) find significant overlap between the representations for number, gender, tense and other morphosyntactic features. Similarly, Stanczak et al. (2022) find overlap in the presentations of morphosyntactic features, particularly for number. Ferrando and Costa-jussà (2024) show the similar structures of circuits for SVA in English and Spanish. The emergence of this kind of interpretability highlights the exciting opportunities presented by MultiBLiMP 1.0 for massively multilingual interpretability.
@@ 3 Linguistic Resources
We briefly describe here our key linguistic resources: Universal Dependencies and UniMorph.
@@ 3.1 Universal Dependencies
Universal Dependencies (UD, v2.15, Zeman et al., 2024) is a multilingual treebank collection, containing 294 treebanks for 167 languages. Besides syntactic dependency relations, UD also contains rich morphological feature annotations that we leverage in our pipeline. An example of a UD-annotated sentence can be found in Fig. 2, step 1.
Treebanks are only included if they consist of the standard variety of their language: dialects and writing standards are split into separate categories, and treebanks of historical or spoken language are discarded. We then consider the union of all remaining treebanks for each language. A list of the excluded treebanks can be found in § A.1.
@@ 3.2 UniMorph
UniMorph (UM, v4.0, Batsuren et al., 2022) is a multilingual collection of morphological feature annotations, which define word-level features such as number, mood, or gender. UM contains annotations for nouns, verbs, and adjectives. It covers 183 languages in total, 81 of which are also covered by UD. Similarly to UD, it uses a universal annotation scheme (Sylak-Glassman et al., 2015), making it highly suitable for our goal of defining a language-independent inflection mechanism to create minimal pairs. UM annotations contain the lemma and morphological features of a word form, making it possible to efficiently disambiguate and create inflections. For example, the English word saw is represented with the following four entries, expressing its various meanings:
lemma  form   featuressee    saw    V;PSTsaw    saw    N;SGsaw    saw    V;NFIN;IMP+SBJVsaw    saw    V;PRS;3;IND;PL Note that features can be left implicit: the first entry of past tense see does not specify person or number, indicating that this form covers all values of those features. We preprocess UM features to be compatible with UD features (SG $\rightarrow$ Sing, PRS $\rightarrow$ Pres, etc.), and transliterate languages that use different scripts from those in UD.
@@ 4 Pipeline
Our pipeline for creating minimal pairs consists of four stages. First, we extract suitable candidates for a phenomenon using dependency parse trees (§4.1). This allows us to determine the key items of interest for a phenomenon, for example, the subject and main finite verb in subject-verb agreement. Next, we validate whether the agreement phenomenon is present in that language based on UD collocation statistics (§4.2). We then inflect one of the words with respect to a particular feature (e.g. number) to create an agreement violation (§4.3). By replacing the item with this inflected form, we create the ungrammatical counterpart for the minimal pair. Finally, we create a minimal pair dataset for each language, balancing the morphological features and amount of pairs (§4.4). We now explain this procedure in more detail; a graphical overview is provided in Figure 2.
@@ 4.1 Candidate Extraction
The 167 UD languages form the initial set of candidate languages we consider for our benchmark. After having excluded some treebanks as described in §3.1, we also set various sentence-level constraints. Specifically, we discard samples containing reparandum dependencies or a Style, Foreign or Typo feature, which typically signal sentences that are malformed in some way.
@@
To extract candidates for a particular linguistic phenomenon, we define filters based on dependency edges, part-of-speech tags and morphological features. In this paper, we focus in particular on various forms of subject-verb and subject-participle agreement. The filters for this phenomenon are defined as follows:
@@
[column sep=1.5em] noun & & aux & verb &
@@
That is, we look for nsubj dependency edges between a noun and verb and collect any auxiliary verb connected to it. For finding subject-participle agreement relations, we place an additional constraint on the morphological features of the verb of VerbForm=Part. We drop subjects containing conjunctions, since these may have conflicting feature values. Furthermore, we drop verbs that have an outgoing expl, csubj:outer, or nsubj:outer dependency, since these indicate clausal subject constructions that may invalidate our minimal pair formation methodology.
@@
PL PL SV 98.9 16.1
PL SG 1.1 83.9
SG SG 98.8 97.5
SG PL 1.2 2.5
PL PL VS 99.0 0.0
PL SG 1.0 100.0
SG SG 99.2 98.2
SG PL 0.8 1.8
Table 1: Dutch and Turkish agreement probabilities for subject-verb number agreement. Significant agreement is denoted in boldface.
@@ 4.2 Agreement Validation
Determining whether a particular phenomenon exists in a given language, and how it can be expressed as a minimal pair, is challenging. We address this with a data-driven procedure that computes the probability of a language containing a phenomenon, based on UD and UM annotations. [^2] An alternative to approaching this from a data-driven perspective would be to use typological databases like Grambank (Skirgård et al., 2023) and WALS (Dryer and Haspelmath, 2013), but their discrete nature and framing agreement in terms of being possible and not compulsory (Baylor et al., 2023) made them unsuitable for our goals. This acts as a final filter on the minimal pairs, ensuring that our perturbed sentences are actually ungrammatical in a given language.
@@
The agreement phenomena we focus on can be turned into a minimal pair by inflecting the agreement feature on one of the words:
@@
Both these constructions are correct, and the construction with the singular verb is even more frequent. A LM provided with this minimal pair would likely assign a higher probability to the sentence with the singular verb and be penalized for it, resulting in an incorrect assessment of the model’s grammatical abilities.
@@
Based on these observations, we define two requirements to determine if a language contains agreement for some feature $\phi$: (1) it must be possible to inflect $\phi$ for both elements of the agreement relation; (2) inflecting $\phi$ for one of the elements must result in a grammatical violation. To test these, we make use of both UD and UM. Req. 1 is easily checked using the morphological features of a word: for all entries of a particular word class in a morphological lexicon, we check if it contains distinct forms for different values of $\phi$. Req. 2 is more challenging.
@@
We approach Req. 2 from a collocational perspective: in a language having strict subject-verb agreement for $\phi$, we expect a very high collocation between nouns and verbs of the same feature value, and a low collocation between nouns and verbs of opposite values. This can be expressed as conditional probability:
$P_{agr}(\phi)=P(\phi^{v}=x|\phi^{n}=x)\approx 1$ (1)
[…]
Equation 1, however, is a simplified approximation of agreement in a language. For instance, French subject-participle agreement depends on the auxiliary: participles inflect for number and gender with an être auxiliary, but not with an avoir auxiliary (elles sont montées vs. elles ont monté). Arabic has flexible subject-verb order and does not inflect for plural in VS order, but only has agreement in SV order for human subjects. Person agreement in Dutch also depends on word order: second person singular verbs in VS order lose their second person ending and coincide with first person (jij doet vs. doe jij). In our pipeline we therefore condition Equation 1 on a binary word order feature, but leave a more fine-grained exploration of agreement conditions open for future work.
@@
Since our procedure depends on the amount of co-occurrence counts we can extract from UD for a language, we need to determine whether the collocation probability is statistically significant. For this, we conduct a binomial test to assess whether Equation 1 was significantly greater than the baseline $p_{0}=0.9$. To rule out agreement, we conduct a binomial test in the opposite direction. If neither test is significant, we mark agreement as being uncertain. For significance we use $\alpha=0.1$ with Bonferroni correction. We only proceed to the minimal pair creation step for the language conditions that have certain and uncertain agreement. The conditions with no agreement are discarded.
Results
Table 1 shows our validation results on subject-verb number agreement for two languages: Dutch, which has number agreement for all conditions, and Turkish, which rarely inflects plural verbs. As expected, Dutch number agreement is marked as significant for both singular and plural, whereas Turkish agreement only passes the test for singular. The number of languages passing the validation for each phenomenon is discussed in §5.

### v2

@@ Abstract
We introduce MultiBLiMP 1.0, a massively multilingual benchmark of linguistic minimal pairs, covering 101 languages and 2 types of subject-verb agreement, containing more than 128,000 minimal pairs. Our minimal pairs are created using a fully automated pipeline, leveraging the large-scale linguistic resources of Universal Dependencies and UniMorph. MultiBLiMP 1.0 evaluates abilities of LLMs at an unprecedented multilingual scale, and highlights the shortcomings of the current state-of-the-art in modelling low-resource languages. [^1] Code: github.com/jumelet/multiblimp
@@ 1 Introduction
Large language models (LLMs) are often trained on highly multilingual corpora, which enable users to interact with them in a wide range of languages (Grattafiori et al., 2024; Üstün et al., 2024). Multilingual evaluation of LLMs, however, has mostly focused on their functional linguistic competence through tasks requiring world knowledge and language understanding (Singh et al., 2024a, b), while little work has assessed their formal linguistic competence, i.e. the “knowledge of rules and statistical regularities of a language” (Mahowald et al., 2024). In multilingual contexts, the latter is commonly approximated intrinsically through perplexity (Chang et al., 2024b) or extrinsically through performance in generative tasks such as translation or summarization (Dang et al., 2024). These approaches do not truly disentangle formal from functional competence. Moreover, they are very coarse-grained and do not inform us on which specific constructions a model does (not) master.
Figure 1: Gemma3-27B accuracy per language on MultiBLiMP 1.0, plotted against language frequency in Common Crawl. Accuracy is measured based on the model assigning a higher probability to a grammatical sentence over a minimally different but ungrammatical sentence. Languages are coloured by their positive or negative deviation from the general trend of accuracy increasing with corpus frequency, highlighting languages that over- or underperform relative to the amount of resources available for them.
Figure 2: Pipeline of the minimal pair creation procedure of MultiBLiMP 1.0
@@
To accelerate progress in this direction, we introduce MultiBLiMP 1.0, a massively multilingual benchmark of linguistic minimal pairs covering two types of subject-verb agreement (subject-finite-verb and subject-participle) for number, person, and gender; created automatically using two large-scale linguistic resources: Universal Dependencies Nivre et al. (2016, 2020); de Marneffe et al. (2021) and UniMorph (Batsuren et al., 2022). MultiBLiMP is not only a benchmark, but also a pipeline for the automatic creation of highly multilingual benchmarks (Fig. 2), which can scale to many more linguistic phenomena. Taking subject-finite-verb [^2] Referred to in the following as subject-verb agreement for simplicity. and subject-participle agreement as a use case, we present a first version of the benchmark including more than 128,000 minimal pairs across 101  languages. We use this to evaluate 42 LLMs, finding that linguistic competence is strongly driven by model size and language frequency in training data (Fig. 1), is acquired during pre-training, and can deteriorate as a result of post-training.
Besides its practical relevance for LLM evaluation, MultiBLiMP enables the study of the learnability of specific linguistic constructions cross-lingually in a unified framework, at a much larger scale than currently possible Mueller et al. (2020), with future applications in LM interpretability (Brinkmann et al., 2025) and quantitative typology (Levshina et al., 2023; Baylor et al., 2024).
@@ 2 Background
2.1 Targeted Syntactic Evaluation
The syntactic abilities of language models have commonly been evaluated using syntactic minimal pairs: sentence pairs that are altered in a minimal way to create a specific grammaticality violation. By comparing a model’s probability judgment of the grammatical sentence to that of the ungrammatical one, we can assess whether it has acquired a notion of the underlying phenomenon (we provide a technical description in §6.1).
@@ 2.1 Targeted Syntactic Evaluation
This approach to test language models was introduced by Linzen et al. (2016), focusing on English subject-verb number agreement. Specifically, they extracted sentences from Wikipedia and flipped the number of the verb without detecting the subject, finding that models at that time struggled considerably with increased distance between the subject and the verb, as well as intervening nouns and other attractors. This work was extended by Marvin and Linzen (2018), who generated sentences with context-free grammars, enabling them to increase the difficulty of agreement across phrases and clauses. BLiMP (Warstadt et al., 2020) extended this approach to a wider range of linguistic phenomena, generating minimal pairs with templates for many linguistic phenomena by sampling from a vocabulary. Hu et al. (2020) and Gauthier et al. (2020) both present similar large-scale collections of syntactic evaluation for English.
Rather than relying on minimal pairs, Pratapa et al. (2021) develop a method to assess the morphosyntactic well-formedness of model outputs based on its dependency parse and a set of rules automatically extracted from parsed corpora. Their approach is complementary to ours as it focuses on the well-formedness of a machine-generate text as opposed to measuring the intrinsic syntactic abilities of a language model on controlled stimuli. However, the requirement of a well-performing parser in each language of interest considerably limits the applicability of their method to a wide set of languages.
Monolingual Non-English Benchmarks
@@
Minimal pair benchmarks have recently been developed for various other languages, using different methods. Some use existing annotated sentences from UD, some craft templates, and some others are built by translating an English dataset. For Chinese, CLiMP (Xiang et al., 2021) generate their minimal pairs with templates adapted from BLiMP, while SLING (Song et al., 2022) edit a constituency treebank with templates and then verify the results with human annotators. ZhoBLiMP (Liu et al., 2024) employ a mixture of adapting paradigms from English BLiMP, and extracting Chinese-specific paradigms from syntax journals and textbooks. JBliMP (Someya and Oseki, 2023) extract minimal pairs from journal articles on Japanese syntax. BLiMP-NL (Suijkerbuijk et al., 2024) create ten sentence pairs per paradigm by hand for Dutch, as well as 90 more sentences generated by ChatGPT and then hand-checked. LINDSEA (Leong et al., 2023) present minimal pairs for Tamil and Indonesian which are created using prompting with GPT3.5 and GPT-4. Kryvosheieva and Levy (2025) create minimal pairs specifically for Swahili noun class agreement, Hindi split ergativity and Basque verb agreement, each chosen for their complexity. Closer to our approach, RuBLiMP (Taktasheva et al., 2024) create pairs from UD-parsed sentences using perturbation rules. Our approach also leverages UD to identify syntactic relations, which we extend multilingually by setting up a language-agnostic minimal pair creation procedure (§4).
@@
Two multilingual minimal pair benchmarks are known to us. The largest, CLAMS (Mueller et al., 2020), creates pairs for English, German, Hebrew, Italian, and Russian for subject-verb agreement by having native speakers translate the sentences from Marvin and Linzen (2018) where applicable. Gulordava et al. (2018) similarly cover English, Hebrew, Italian, and Russian for various types of agreement using sentences from UD treebanks, in which content words are swapped out with others of the same part-of-speech to test the models’ reliance on semantics. Two other notable related efforts present a multilingual benchmark for linguistic acceptability—for which a model is fine-tuned to predict the grammaticality of a single sentence instead of judging a pair (Warstadt et al., 2019)—ScaLA, which is part of ScandEval (Nielsen, 2023), covering the 5 Scandinavian languages, and MELA (Zhang et al., 2024), covering a diverse of 10 languages.
We see MultiBLiMP 1.0 as complementary to all these efforts. While language-specific benchmarks can provide tremendous depth on the syntactic system of a language, it will prove challenging to scale such efforts to a wide variety of languages. Our approach, on the other hand, focuses on ‘wide coverage’ and can therefore be of use for both the evaluation of highly multilingual LLMs and for quantitative typological studies.
@@ 2.2 Representation Sharing in LLMs
Various works have evaluated the degree to which multilingual models make use of shared syntactic representations across languages. Early studies analyzed mBERT (Devlin et al., 2019) and smaller-scale LSTM models trained on a few languages, finding mixed evidence of sharing (Pires et al., 2019; Mueller et al., 2020; Dhar and Bisazza, 2021; Chi et al., 2020; Choenni and Shutova, 2022). Focusing on modern-scale LLMs, Wendler et al. (2024) find that models trained on English-dominated corpora use English to some degree as an internal ‘pivot language’ to solve tasks in other languages. Brinkmann et al. (2025) find significant overlap between the representations for number, gender, tense and other morphosyntactic features. Similarly, Stanczak et al. (2022) find overlap in the presentations of morphosyntactic features, particularly for grammatical number. Ferrando and Costa-jussà (2024) show the similar structures of circuits for subject-verb agreement in English and Spanish. The emergence of this kind of interpretability findings highlights the exciting opportunities presented by MultiBLiMP 1.0 for massively multilingual interpretability.
@@ 3 Linguistic Resources
We briefly describe here the key linguistic resources that we use to create MultiBLiMP: Universal Dependencies and UniMorph.
@@ 3.1 Universal Dependencies
Universal Dependencies Nivre et al. (2016, 2020); de Marneffe et al. (2021) is a multilingual treebank collection, containing 296 treebanks for 168 languages. Besides syntactic dependency relations, UD also contains rich morphological feature annotations that we leverage in our pipeline. An example of a UD-annotated sentence can be found in Fig. 2, step 1. We use UD version 2.15 for the 1.0 release of MultiBLiMP Zeman et al. (2024).
We exclude treebanks of spoken language, historical varieties, as well as certain genres, e.g. technical manuals. Generally, we only exclude treebanks for low-resource languages where absolutely necessary, while for high-resource languages, we also exclude very small treebanks that will be of little additional benefit, but may introduce annotation inconsistencies. A full list of the excluded treebanks with reasons for exclusion can be found in § A.1. We then consider the union of all remaining treebanks for each language.
@@ 3.2 UniMorph
UniMorph (UM, v4.0, Batsuren et al., 2022) is a multilingual collection of morphological feature annotations, which define word-level features such as number, mood, or gender for nouns, verbs, and adjectives. It covers 183 languages in total, 81 of which are also covered by UD. Similarly to UD, it uses a universal set of features (Sylak-Glassman et al., 2015), making it highly suitable for our goal of defining a language-independent inflection mechanism to create minimal pairs. UM entries contain the lemma and morphological features of a word form, making it possible to efficiently disambiguate and create inflections. For example, the English word saw is represented with the following four entries, expressing its various meanings:
lemma  form   featuressee    saw    V;PSTsaw    saw    N;SGsaw    saw    V;NFIN;IMP+SBJVsaw    saw    V;PRS;3;IND;PL Note that features can be left implicit: the first entry of past tense see does not specify person or number, indicating that this form covers all values of those features. We preprocess UM features to be compatible with UD feature values (SG $\rightarrow$ Sing, PRS $\rightarrow$ Pres, etc.), and transliterate languages that use different scripts from those in UD.
@@ 4 Pipeline
Our pipeline for creating minimal pairs consists of four stages. First, we extract suitable candidates for a phenomenon using dependency parse trees (§4.1). This allows us to determine the key items of interest for a phenomenon, for example, the subject and verb in subject-verb agreement. Next, we validate whether the agreement phenomenon is present in that language based on UD collocation statistics (§4.2). We then inflect one of the words with respect to a particular feature (e.g. number) to create an agreement violation (§4.3). By replacing the item with this inflected form, we create the ungrammatical counterpart for the minimal pair. Finally, we create a minimal pair dataset for each language, balancing the morphological features and number of pairs (§4.4). We now explain this procedure in more detail; a graphical overview is provided in Figure 2.
@@ 4.1 Candidate Extraction
The 167 UD languages form the initial set of candidate languages we consider for our benchmark. After having excluded some treebanks as described in §3.1, we also set various sentence-level constraints. Specifically, we discard sentences containing reparandum dependencies or a Style, Foreign or Typo feature, which typically signal sentences that are malformed in some way.
@@
To extract candidates for a particular linguistic phenomenon, we define filters based on dependency edges, part-of-speech tags and morphological features. In this paper, we restrict our attention to various forms of subject-verb and subject-participle agreement. The filters for this phenomenon are defined as follows:
@@
[column sep=1.5em] subject & & aux & verb &
@@
That is, we look for nsubj dependency edges between a noun and a verb and collect any auxiliary verb connected to it. If the verb or one of its auxiliaries has the feature VerbForm=Finite, it is considered as a candidate for subject-verb agreement. For finding subject-participle agreement relations, we place an additional constraint on the morphological features of the verb of VerbForm=Part. We drop subjects containing conjunctions, since these may have conflicting feature values. Furthermore, we drop verbs that have an outgoing expl, csubj:outer, or nsubj:outer dependency, since these indicate clausal subject constructions that may invalidate our minimal pair formation methodology.
@@
PL PL SV 0.989 0.161
PL SG 0.011 0.839
SG SG 0.988 0.975
SG PL 0.012 0.025
PL PL VS 0.990 0.000
PL SG 0.010 1.000
SG SG 0.992 0.982
SG PL 0.008 0.018
Table 1: Dutch and Turkish agreement probabilities for subject-verb number agreement. Significant agreement is denoted in boldface. WO stands for word order, either subject verb (SV) or verb subject (VS).
@@ 4.2 Agreement Validation
Determining whether a particular phenomenon exists in a given language, and how it can be expressed as a minimal pair, is challenging. We address this with a data-driven procedure that computes the probability of a language containing a phenomenon, based on UD and UM annotations. [^4] An alternative to approaching this from a data-driven perspective would be to use typological databases like Grambank (Skirgård et al., 2023) and WALS (Dryer and Haspelmath, 2013), but their discrete nature and framing agreement in terms of being possible and not compulsory (Baylor et al., 2023) make them unsuitable for our goals. This acts as a final filter on the minimal pairs, ensuring that our perturbed sentences are actually ungrammatical in a given language. [^5] In addition to this data-driven procedure, we also conduct extensive spot checks to ensure our procedure yields valid minimal pairs.
@@
The agreement phenomena we focus on can be turned into a minimal pair by changing the agreement feature on a word and re-inflecting it:
@@
Both these constructions are correct, and the construction with the singular verb is even more frequent. A model provided with this minimal pair would likely assign a higher probability to the sentence with the singular verb and be penalized for it, resulting in an incorrect assessment of the model’s grammatical abilities.
@@
Based on these observations, we define two requirements to determine if a language contains agreement for some feature $\phi$ (Baker, 2008):
  1. 1.
    $\phi$ must be specified for both elements of the agreement relation, either lexically inherent (la fille est tombée; the girl fell) or morphologically realized (les garçons sont tombés; the boys fell). This way, for example, we rule out subject-verb gender agreement in English, since gender is not specified on verbs. This is easily validated using UniMorph and UD by checking whether both nouns and verbs are annotated for the feature.
  2. 2.
    Re-inflecting one of the elements with a different value for $\phi$ for one of the elements must result in a grammatical violation. This requirement is more challenging to test.
@@
We approach Requirement 2 from a collocational perspective: in a language having strict subject-verb agreement for $\phi$, we expect very high co-occurrence between nouns and verbs of the same feature value, and low co-occurrence between nouns and verbs of contrasting values. This can be expressed as a conditional probability:
$P_{agr}(\phi^{v}=x|\phi^{n}=x,\textup{WO})\approx 1$ (1)
where $\phi^{v}$ and $\phi^{n}$ denote the feature of verb and noun, $x$ a specific feature value (e.g. sg or pl), and WO the subject-verb word order (SV/VS). If this probability is close to 1, it implies that the feature of the noun co-occurs highly with that of the verb, and the probability of the verb having a contrasting feature value is close to 0.
Agreement can depend on various extra factors, such as word order. For example, person agreement in Dutch depends on this: second person singular verbs in VS order lose their second person ending and coincide with first person (jij doet vs. doe jij). A different example is French, for which subject-participle agreement depends on the auxiliary: participles inflect for number and gender with an être auxiliary, but not with an avoir auxiliary (elles sont montées vs. elles ont monté). Arabic has flexible subject-verb order and does not inflect for plural in VS order, but only has agreement in SV order for human subjects. In our pipeline we therefore condition Equation 1 on a binary word order feature, but leave a more fine-grained exploration of agreement conditions open for future work, for instance by extending the data-driven agreement extraction approach of Chaudhary et al. (2020) with word order and lexical features.
Agreement Conditions
We compute agreement probabilities for each language, verb type (finite main verb, finite auxiliary or participle), feature value, and word order. We refer to this configuration as an agreement condition. Probabilities for each agreement condition are computed by counting feature co-occurrences based on the morphological annotations in the UD tree, using the dependency relations we extract from UD (§4.1). In case a feature is not annotated, we look it up in UniMorph and, if present, take the value from there. We only count co-occurrences for cases where a feature is defined for both the subject and the verb (e.g. number is not specified for past tense verbs in English). Based on the co-occurrence counts of the feature values, we then compute the conditional probability distribution of Equation 1.
Note that it is possible this way for a language to have strict agreement in only one direction. We provide an example of this in Table 1, in which we provide subject-verb number agreement for two languages: Dutch, which has strict number agreement for all conditions, and Turkish, which does not require number agreement for plural subjects. Dutch number agreement is marked as significant for both singular and plural, whereas Turkish agreement only demonstrates agreement for singular items.
@@
Since our procedure depends on the amount of co-occurrence counts we can extract from UD for a language, we need to determine whether the collocation probability is statistically significant. For this, we conduct a binomial test to assess whether Equation 1 was significantly greater than the baseline $p_{0}=0.9$. To rule out agreement, we conduct a binomial test in the opposite direction. If neither test is significant, we mark agreement as being uncertain. For significance we use $\alpha=0.1$ with Bonferroni correction. We only proceed to the minimal pair creation step for the agreement conditions that have certain and uncertain agreement. The conditions with no agreement are discarded.
Figure 3: Number of minimal pairs per language in MultiBLiMP, split out for certain and uncertain agreement cases using the agreement detection procedure of §4.2. Note the log-scale on the y-axis.